# SpecTra: Enhancing the Code Translation Ability of Language Models by Generating Multi-Modal Specifications

### v2

@@ 4.1 Evaluating the Generated Specifications
[…]
@@
[…]
4.2 Impact of Individual Specification Modalities
First, we evaluate the efficacy of each specification modality individually. For each of the 300 C and JavaScript programs, we attempt to perform C to Rust, C to Go, and JavaScript to TypeScript translation using gpt-4o. While generating translations, we provide each specification modality individually along with the program and obtain one candidate translation per modality.
Our hypothesis was that different modalities would act in a complementary fashion, whereby each modality might provide some extra information for translation that the other modalities do not provide. To investigate this, we evaluate all the translations and plot a Venn Diagram in Figure 4 to visualize the contributions of each individual specification modality. While there exists some overlap, the different modalities are indeed complementary to one another.
4.3 Impact of Multiple Specification Modalities
To evaluate the impact of multiple modalities, we measure the number of accurate translations produced SpecTra (at various steps) against the baseline (where the prompt contains no additional specs) at various pass@k. Note that each pass at k corresponds to the equivalent step-k, i.e., pass@1 corresponds to step 1, and so on. We also measure % improvement over the baseline at various steps/pass@k’s. In this comparison, a positive percentage score indicates that, compared to the baseline, more source programs are accurately translated by SpecTra. Our findings are summarized in Sections 4.1 and • ‣ 4.3. The tabulated findings may be summarized as below:
SpecTra produces more correct translations compared to baselines: SpecTra generally shows improvement over the baseline. The largest relative gains are on the C to Rust translation task, and in particular, Gemini shows relative improvements of between 19% and 23%. The best-performing base models are GPT4o and Claude, but they too are able to benefit from using SpecTra.
The largest relative increases generally occur for models that have the lowest baseline performances: For example, Deepseek and Gemini are the two worst models at the C to Rust task, and their “improvement %” of up to 23% are the highest of all models on this task. Conversely, Deepseek’s baseline performance is much better on C to Go, and its performance gain with SpecTra is low. However, Granite is an exception to this trend. Its baseline performance is poor, but it is unable to effectively use our specifications for any of the tasks.
[…]
@@ 4.3 Impact of Multiple Specification Modalities
[…]
step3 vs. step1
step2 vs. step1
step3 vs. step1
step2 vs. step1
step3 vs. step1
[…]
@@
SpecTra also helps improve Javascript to Typescript translation albeit with considerably less pronounced gains compared to the other two translation tasks: Translation from Javascript to Typescript (• ‣ Section 4.3) indicate an similar trend the other two translation pairs in that SpecTra generally improves results over the baseline, with notable percentage improvements for GPT4o and Gemini. However, overall improvements are far less pronounced compared to other translation tasks. We conjecture that this is due to the high degree of semantic and syntactic similarity between the two languages. JavaScript and TypeScript share many features and structures, and the translation process is inherently simpler and less reliant on detailed specifications to achieve correctness. This similarity reduces the need for extensive specification guidance, as the language model can more easily infer the correct translations based on existing structural and semantic parallels. Further, providing additional specifications in fact can be detrimental to some smaller models like Granite and Deepseek. For instance, the step-1 accuracy for Granite on JavaScript to TypeScript translation is 15% worse than the baseline model without specifications.
Lastly, in order to assess our hypothesis that providing additional modalities of specification when the the static specification in insufficient for correct translation can benefit translation, we compare the following in • ‣ Section 4.3:
The percentage (%) additional correct translation for SpecTra by using either 2 or 3 different specification modalities individually (i.e., step 2 and step 3 in • ‣ Section 4.3) compared to only one single specification modality (step-1).
The percentage (%) additional correct translation for baseline model at Pass@3 and Pass@2 versus the same baseline model at Pass@1. We note that the baseline uses no specifications for translation.
The findings summarized in • ‣ Section 4.3 demonstrate that incorporating additional specification modalities when static specifications are inadequate can significantly enhance translation accuracy. In all the cases, additional steps increase the number of correct translations (as evidenced by positive and higher percentages shown in blue). We also note that Step 3 vs. Step 1 is greater than Step 2 vs. Step 1, indicating that even informal descriptions can help increase the number of correct translations. Lastly, it is worth noting that although there are marginal (albeit positive) improvements for Javascript to Typescript, models with a lower baseline accuracy like Granite see considerable improvements with providing additional specification modalities - a 22% improvement when using up to two kinds of specifications and 26% when using up to 3 kinds of specifications.
5 Related Work
5.1 Code Translation with LLMs
@@
Roziere et al. (2020) proposed TransCoder, an early attempt to tackle code translation using transformer models. They proposed Back Translation, a fully unsupervised approach. Ahmad et al. (2022) refined Back Translation with the addition of a code summarization task. In contrast to unsupervised learning, supervised learning requires aligned parallel data across pairs of programming languages, which is challenging to obtain at scale. Roziere et al. (2021) address this challenge by generating synthetic aligned data using another language model and filtering out incorrect translations using generated unit tests. The last few years have seen the development of large foundation models for code and natural language, that are pretrained on massive amounts of data with an autoregressive objective function. These can perform a variety of code-related tasks, including translation. Pan et al. (2024) and Yang et al. (2024) perform thorough empirical evaluations of the translation ability of these models.
5.2 Generating Program Specifications with LLMs
@@
The three kinds of program specifications that we consider in this paper are test cases, natural language descriptions, and static specifications. There have been a few recent papers that use LLMs to generate test cases for competitive coding problems. EvalPlus Liu et al. (2023) uses ChatGPT to generate test cases, and mutates them in order to rigorously test LLM-generated code solutions. CodeT Chen et al. (2022) uses LLM-generated test cases to filter out code solutions from a list of LLM-generated candidate solutions. Recently, Yang et al. (2024) proposed to generate test cases, provide them along with the code to an LLM for translation, and use test feedback to iteratively repair the generated translation.
Separately, there has been a long line of work on using transformer models for generating natural language descriptions of code Ahmad et al. (2022); Gao et al. (2023); Gong et al. (2022); Gao and Lyu (2022); Wu et al. (2021); Tang et al. (2021). There has been comparatively less research into generating static specifications. Endres et al. (2023) transform natural language descriptions of functions into assert statements and use these assertions to detect incorrect code.
Parsel Zelikman et al. (2023) is a recent approach that generates and utilizes multiple kinds of specifications. Starting with a problem description, Parsel first generates a program sketch consisting of function-level descriptions and test cases, and then generates code to complete the sketch.
6 Conclusion
@@
In this paper, we have presented SpecTra, an approach to 1) generate self-consistent multi-modal specifications, and 2) use these specifications to improve the program translation performance of large language models. We show relative performance gains of up to 26% across multiple models and languages pairs. Our results indicate the possibility for further research into generating high-quality program specifications as a cheap and efficient way to improve the performance of real-world LLMs on code translation. Future work could explore generating specifications in formal language or assert statements, which can then be automatically cross-verified against the source code. Another promising direction is to use these specifications for downstream tasks other than translation.
7 Ethical Impact and Potential Risks
The inherent risk associated with using LLMs is that the code they produce is not guaranteed to be correct. Even though the generated code may pass some given test cases, there may be other corner cases and unexpected behavior that are not exposed. If code translation with LLMs starts being used at a large scale without adequate testing, it could create an ecosystem of untested, unproven, potentially insecure code. Although our paper generates different kinds of specifications, it does not cross-verify the final translations against these specifications. We encourage the development of specification generation and formal verification methods for LLM-generated programs as a means to counter potential security vulnerabilities.
8 Limitations
One limitation of our approach for generating static specifications is that these specifications are in natural language and cannot be precisely validated against the source program. A self-consistent static specification, while likely to be correct, is not guaranteed to be correct.
Our paper deals only with code translation in a competitive coding setting, where we have relatively simple standalone files which take text input from STDIN and write text to STDOUT. However, real-world legacy code is far more complex, often involving multiple files with hundreds of lines of code each, and complex inter-dependencies. Testing such programs is also non-trivial, and may involve, say, simulating a server and querying it with HTTP requests, or setting up a mock database to simulate reads and writes. Performing code translation in this kind of setting is an important challenge and a natural direction to extend our work.
Appendix A Example Prompts
Figure 5: The first step of our approach involves generating multiple candidate specifications from a given program. This is the actual output of gpt-4o on our motivating example from Figure 1. Figure 6: Filtering out incorrect specifications. We re-generate C code using each specification, and pick the specifications corresponding to re-generated C programs that pass the test case. Figure 7: Using different modalities of specifications to perform specification-augmented translation. These are the actual prompts and specifications generated by SpecTra. - •

### v3

@@ 5. Empirical Results
| Baseline | 176 | 192 | 202 | ||||
| Gpt4o | SpecTra | 196 | 219 | 220 | 11% | 14% | 9% |
| Baseline | 112 | 125 | 134 | ||||
| Gpt4o-mini | SpecTra | 124 | 150 | 156 | 11% | 20% | 16% |
| Baseline | 180 | 196 | 201 | ||||
| Claude | SpecTra | 208 | 226 | 227 | 16% | 15% | 13% |
| Baseline | 160 | 174 | 181 | ||||
| Gemini | SpecTra | 171 | 194 | 199 | 7% | 12% | 10% |
| Baseline | 95 | 106 | 117 | ||||
| Llama | SpecTra | 106 | 130 | 135 | 12% | 23% | 15% |
| Baseline | 101 | 117 | 124 | ||||
| Mistral | SpecTra | 147 | 159 | 162 | 46% | 36% | 31% |
| C to Go | |||||||
| No. correct | improvement % | ||||||
@@
| Baseline | 183 | 207 | 218 | ||||
| Gpt4o | SpecTra | 204 | 228 | 233 | 11% | 10% | 7% |
| Baseline | 143 | 158 | 167 | ||||
| Gpt4o-mini | SpecTra | 158 | 177 | 179 | 10% | 12% | 7% |
| Baseline | 238 | 246 | 250 | ||||
| Claude | SpecTra | 243 | 262 | 267 | 2% | 6% | 7% |
| Baseline | 161 | 177 | 184 | ||||
| Gemini | SpecTra | 176 | 198 | 203 | 9% | 12% | 10% |
| Baseline | 121 | 149 | 165 | ||||
| Llama | SpecTra | 134 | 168 | 176 | 11% | 13% | 7% |
| Baseline | 126 | 152 | 166 | ||||
| Mistral | SpecTra | 159 | 176 | 179 | 26% | 16% | 8% |
RQ1. Overall Performance
In this section, we evaluate SpecTra against a baseline that does not use specifications.
Evaluation: We measure the number of accurate translations produced by SpecTra (at various steps) against the baseline (where the prompt contains no additional specs) at various pass@k. Note that each pass at k corresponds to the equivalent step-k, i.e., pass@1 corresponds to step 1, and so on. We also measure % improvement over the baseline at various steps/pass@k’s. In this comparison, a positive percentage score indicates that, compared to the baseline, more source programs are accurately translated by SpecTra. Our findings are summarized in Tables˜2 and 3.
Discussion: The tabulated findings may be summarized as below:
SpecTra produces more correct translations compared to baselines: SpecTra generally shows improvement over the baseline. The largest relative gains are on the C to Rust translation task, and in particular, Mistral shows relative improvements of between 31% and 46%. The best-performing base models are GPT4o and Claude, but they too are able to benefit from using SpecTra.
The largest relative increases generally occur for models that have the lowest baseline performances, and vice-versa: For example, Mistral and Llama are the two worst models at the C to Rust task, and the highest “improvement %” on this task is on Mistral (31-46%) followed by Llama (23%). The converse is also true - Claude is the best performing model on C to Go, and its performance gain (2-7%) is the lowest of all models on this task.
Table 3. Comparing SpecTra to the baseline for Javascript to Typescript. The best absolute numbers for each column are in orange, percentage increases are in blue, percentage decreases are in pink. SpecTra does improve performance, but gains are limited. No. correct improvement % pass@1
@@
Baseline 289 295 295 Gpt4o SpecTra 284 292 294 -2% -1% -0.3% Baseline 282 284 285 Gpt4o-mini SpecTra 279 289 289 -1% 2% 1% Baseline 282 285 286 Claude SpecTra 283 292 292 0.4% 2% 2% Baseline 284 289 290 Gemini SpecTra 282 291 292 -1% 1% 1% Baseline 268 278 283 Llama SpecTra 270 284 285 1% 2% 1% Baseline 285 289 289 Mistral SpecTra 280 291 292 -2% 1% 1% -
@@
SpecTra also helps improve Javascript to Typescript translation albeit with considerably less pronounced gains compared to the other two translation tasks: SpecTra generally improves results over the baseline (Table˜3), with consistent percentage improvements for Claude and Llama. However, overall improvements are far less pronounced compared to other translation tasks. We conjecture that this is due to the high degree of semantic and syntactic similarity between the two languages. JavaScript and TypeScript share many features and structures, and the translation process is less reliant on detailed specifications to achieve correctness. The LLM can infer the correct translations based on existing structural and semantic parallels. We also see in Table˜3 that specifications can be detrimental in certain cases. On manual inspection, we find that the JavaScript specifications are extremely verbose 444Since JavaScript does not require procedural code to be enclosed within functions, we ask the LLM to partition the code into logically coherent segments. In practice, this leads to as many as 20 segments in some programs, each with their own preconditions and postconditions., and add a lot of tokens to the input. As the length of the context increases, LLMs are more prone to hallucinations and other errors (Peng et al., 2023). Future work could explore generating concise specifications for JavaScript.
RQ2. Impact of Individual Specification Modalities
Here we evaluate the efficacy of each specification modality individually. Our hypothesis is that different modalities would act in a complementary fashion, whereby each modality might provide some extra information for translation that the other modalities do not provide.
Evaluation: For each of the 300 C and JavaScript programs, we attempt to perform C to Rust, C to Go, and JavaScript to TypeScript translation using gpt-4o. While generating translations, we provide each specification modality individually along with the program and obtain one candidate translation per modality.
Discussion: We evaluate all the translations and plot a Venn Diagram in Figure˜2 to visualize the contributions of each individual specification modality. Because of a lack of space, we show only the C to Rust diagram here; the others are in the appendix. There are 71 programs that are solved with all three modalities, showing that there is a substantial overlap. However, the remaining 213 programs are solved with some combination of two or fewer modalities, showing that the different modalities are indeed complementary.
RQ3. Impact of Design Choices
In this RQ, we first investigate the effectiveness of our multi-stage approach in SpecTra, where we provide each of the modalities of specifications one by one, as opposed to providing all of them together. Next, we evaluate the impact of the self-consistency check on translation quality.
Evaluation 1: We take a single setting - GPT-4o for C to Rust translation, and compare SpecTra with a baseline where we provide all the specifications together in the same prompt, and take the top candidates to compute pass@k. The results are shown in Table˜4(a).
Discussion 1: We can see that SpecTra is better than the combined baseline. This justifies our intuition in Section˜3.3 that using all modalities simultaneously can create a surfeit of information, expanding the size of the LLM input and degrading its performance.
Evaluation 2: We again take GPT-4o for C to Rust translation, and compare SpecTra with a baseline where we take the first specification generated by our specification generation model, with no correctness check. We call this a “one-shot specification”. The results are shown in Table˜4(b).
Discussion 2: The self-consistency check makes a noticeable positive difference for translation using Static and I/O specifications. In fact, for I/O specifications, translation with one-shot specifications is worse than translation with no specifications. However, translation with Descriptions is unaffected by the self-consistency check. We hypothesize that this is because of the imprecise nature of natural language. The inherent ambiguity of the specification could mask small semantic mistakes. Likewise, this same ambiguity could be a threat to the validity of the self-consistency filter for descriptions.
pass@1 |
pass@2 |
pass@3 |
pass@1 |
pass@2 |
pass@3 |
| w/o Specifications | 176 | 192 | 202 | |||||||||
| All Specs Together | 195 | 215 | 219 | 11% | 12% | 8% | ||||||
| SpecTra | 196 | 219 | 220 | 11% | 14% | 9% | ||||||
Static |
I/O |
Desc |
Static |
I/O |
Desc |
| # Programs w. Spec | 189 | 294 | 134 | |||||||||||
| w/o Specifications | 131 | 174 | 88 | |||||||||||
| One-shot Spec | 140 | 170 | 97 | 7% | -2% | 10% | ||||||||
| Self-consistent Spec | 146 | 188 | 96 | 11% | 8% | 9% | ||||||||
RQ4. Evaluating the Generated Specifications
We would like to determine whether our specification generation process is able to generate high-quality self-consistent specifications.
Evaluation: To do this, we check how many programs in our benchmark dataset we are able to generate self-consistent specifications for. For each program, we generate up to candidate tests, and up to static specifications and descriptions. In all the cases, we stop as soon as we find a self-consistent specification. We use gpt-4o as our LLM for both the “forward” specification generation as well as the “reverse” code generation process. We also measured the generated test coverage for C programs using the gcov tool.
| Lang | Total | Static | I/O | Desc |
| C | 300 | 189 (63.0%) | 294 (98.0%) | 134 (44.7%) |
| JS | 300 | 199 (66.3%) | 296 (98.7%) | 189 (63.0%) |
Discussion: The number of self-consistent specifications generated by gpt-4o of each category is shown in Table˜5. The average coverage was 91.5%, which is an indicator that the tests are of high quality. For the self-consistent descriptions and static specifications, the fact that we are able to recover a functioning program from them is a strong validation of their quality. We also manually inspected several of them and confirmed that they are a good representation of the program’s behavior.
6. Case Study of Translating Projects
6.1 Approach
We would like to adapt our specification-driven approach to translate full software projects. This setting presents multiple challenges:
The limitation of the LLM context window. The entire project may consist of thousands of lines of code, so it might not fit into an LLM input context window.
Solution: We need to decompose the project into smaller pieces that can be individually translated. In the competitive coding setting, our Static and NL specifications were generated at the function level, so here too, we generate specifications at the granularity of each function.
How to evaluate the correctness of each individual function translation. If a function is translated from one language to another, the code may no longer compile because of other functions that call this function. Further, although we may have end-to-end test cases, it is difficult to evaluate a standalone function translation.
Solution: Our solution is to use cross-language foreign function APIs to enable functions in one language to call functions in another language, and vice-versa. That way, the entire code still compiles and runs end-to-end. The translated function becomes a modified link in a chain of functions, and if we run end-to-end tests, we can see whether the modified function preserves the overall program’s correctness with respect to these tests.
The kinds of errors that LLMs make while translating complex functions with many dependencies. There are inevitably some errors due to missing imports, redeclaring variables, and other “shallow” compilation errors that are not due to lack of understanding of function semantics.
Solution: We perform a finite number of iterations of compiler feedback-aided repair, where we take the compiler log and give it to the LLM to re-generate its solution. We do this both for the baseline as well as for SpecTra.
How to curate different kinds of specifications in this setting. Generating specifications for individual functions using an LLM is easy, but validating these using self-consistency requires a different approach to that presented in Section˜3.2.
Solution: To generate static specifications for a single function, we ask the LLM to generate a precondition and postcondition. Then, using this precondition and postcondition, we generate a back-translation as in Section˜3.2, but only of this single function. We swap out the original version and swap in the back-translated version, and run end-to-end tests to verify the correctness of the back-translation.
We could follow a similar process as above for NL specifications, but in the project we chose, most functions were already annotated with a docstring that briefly described their behavior. We extracted these docstrings and used them as our NL specifications.
When it comes to I/O specifications, program-level or end-to-end test cases are unlikely to be of much assistance to the LLM in a full project setting. If we are translating a single function in the middle of a chain of function calls, then it is extremely unlikely that end-to-end tests can provide information to aid in translating this single function. Our solution is to instrument the entire program to log the arguments and return values of each function. We run the provided end-to-end tests and collect these intermediate values, which we then provide to the LLM as I/O specifications. For this paper, we use the provided end-to-end tests, but future work could investigate using fuzzing to generate tests.
6.2 Study Results
Evaluation: We run SpecTra with suitable modifications as described in Section˜6.1 on our dataset consisting of 24 functions from cat. We first generate specifications for each individual function, and then use these specifications to translate each function. After each translation, we verify correctness by running end-to-end tests. We measure pass@k in the same manner as in Section˜4.3, except for each individual function instead of each program. We also allow up to 3 rounds of compiler-feedback based LLM debugging before evaluating each candidate translation.
Discussion: We are able to generate self-consistent static specifications for 10 out of 24 functions. 18 out of 24 functions have docstrings, and all the functions have input-output specifications collected through instrumentation. Table˜7 shows the results of using these specifications to translate the 24 functions. We can see that SpecTra yields a small improvement in pass@k for all values of k. Although this is a small delta, we expect that this will translate into a larger gain in terms of absolute results if we evaluate our technique on a larger dataset. Additionally, as we observed for standalone programs, using multiple specification modalities together is not as effective as our multi-stage pipeline. In fact, this actually degrades performance slightly compared to a baseline without specifications.
Table˜6 shows an example of a function in C translated to Rust, along with its static and I/O specifications. We can see that the generated specifications are semantically meaningful, and provide information that SpecTra is able to effectively use the information from these specifications to guide its translations.
pass@1 |
pass@2 |
pass@3 |
pass@1 |
pass@2 |
pass@3 |
| w/o Specifications | 10 | 11 | 12 | |||||||||
| All Specs Together | 8 | 11 | 11 | -20% | 0% | -8% | ||||||
| SpecTra | 11 | 12 | 13 | 10% | 9% | 8% | ||||||
| Total #Func. | 24 | |||||||||||
7. Related Work
@@
Code Translation with LLMs. Roziere et al. (2020) proposed Trans-Coder, an early attempt to tackle code translation using transformer models. They proposed Back Translation, a fully unsupervised approach. (Ahmad et al., 2022) refined Back Translation with the addition of a code summarization task. In contrast to unsupervised learning, supervised learning requires aligned parallel data across pairs of programming languages, which is challenging to obtain at scale. (Roziere et al., 2021) address this challenge by generating synthetic aligned data using another language model and filtering out incorrect translations using generated unit tests. The last few years have seen the development of large foundation models for code and natural language, that are pretrained on massive amounts of data with an autoregressive objective function. Pan et al. (2024) and Yang et al. (2024) perform thorough empirical evaluations of the translation ability of these models. Recently, there has been a lot of interest in translating full software repositories (Shiraishi and Shinagawa, 2024; Ibrahimzada et al., 2024; Shetty et al., 2024; Zhang et al., 2024; Nitin et al., 2025). Although all of these works differ in their approach and setting, the only one that utilizes any kind of specifications is Shetty et al. (2024). They use carefully crafted Rust-specific properties collected through dynamic analysis to guide the LLM. Our specifications are much broader in scope, however.
@@
Generating Program Specifications with LLMs. The three kinds of program specifications that we consider in this paper are test cases, natural language descriptions, and static specifications. There have been a few recent papers that use LLMs to generate test cases for competitive coding problems. EvalPlus (Liu et al., 2023) uses ChatGPT to generate test cases, and mutates them in order to rigorously test LLM-generated code solutions. CodeT (Chen et al., 2022) uses LLM-generated test cases to filter out code solutions from a list of LLM-generated candidate solutions. Yang et al. (2024) proposed to generate test cases, provide them along with the code to an LLM for translation, and use test feedback to iteratively repair the generated translation. Separately, there has been a long line of work on using transformer models for generating natural language descriptions of code (Ahmad et al., 2022; Gao et al., 2023; Gong et al., 2022; Gao and Lyu, 2022; Wu et al., 2021; Tang et al., 2021). Tang et al. ([n. d.]) and Saha et al. (2024) both propose to generate natural language explanations as an intermediate step for program translation. There has been comparatively less research into generating static specifications. Endres et al. (2023) transform natural language descriptions of functions into assertions and use them to detect incorrect code. Parsel (Zelikman et al., 2023) first generates a program sketch consisting of function-level descriptions and test cases, and then generates code to complete the sketch. Park et al. (2024) explore synthesizing formal static specifications for nondeterministic programs.
8. Threats to Validity
Our results have several threats to validity.
Static Specs expressed in NL. Our static specs are written in natural language, making it difficult to validate them precisely against the source program. While a self-consistent specification is a strong indication of correctness, it does not guarantee correctness. The absence of formal verification introduces potential ambiguities or misinterpretations. One way to address this issue would be to use formal specifications. However, we opted against this approach for two primary reasons:
The C programming language lacks a succinct and expressive mechanism for specifying complex logical assertions, such as quantification over data structures (e.g., “for all” conditions).
Our preconditions and postconditions are written in NL, aligning with how LLMs are trained and optimized mostly on text corpora.
Computational Overhead of SpecTra SpecTra has computational overhead over a single-step translation using LLMs. Even in the best case, where we find a self-consistent specification in the very first step, translating a program involves at least 2 additional LLM calls. Thus, there is a tradeoff between better performance (SpecTra) and fewer LLM calls (vanilla translation). However, once specifications have been generated for a single source program, they can be used for multiple downstream tasks without any additional overhead.
Dataset Size and Statistical Significance Our dataset for full-project translation is relatively small, limiting the statistical significance of our results in this setting. While our findings suggest that our approach can be extended to full-project translation, they should be interpreted as preliminary evidence rather than definitive proof of scalability. Nonetheless, we present this as a proof of concept, demonstrating that our method is viable in a full-project setting. Future work will involve scaling up our dataset and conducting more extensive evaluations to assess generalizability.
9. Conclusion
@@
In this paper, we have presented SpecTra, an approach to 1) generate self-consistent multi-modal specifications, and 2) use these specifications to improve the program translation performance of large language models. We show relative performance gains of up to 46% across multiple models and languages pairs. Our results indicate the possibility for further research into generating high-quality program specifications as a promising way to improve the performance of real-world LLMs on code translation. Future work could explore generating specifications in formal language or assert statements, which can then be automatically cross-verified against the source code. Another promising direction is to use these specifications for downstream tasks other than translation.
Appendix A Additional Diagrams